A multiple choice question on an exam has four possible responses—(A), (B), (C), and (D). When 440 students take the exam, 132 give response (A), 88 give response (B), 176 give response (C), and 44 give response (D). Write out the frequency distribution, relative frequency distribution, and percent frequency distribution for these responses. (Round your relative frequency answers to 1 decimal place. Omit the "%" sign in your response.) Frequency Distributions for Student Responses Category / Class Frequency Relative Frequency Percent Frequency A % B % C % D %
Answer #1
Response Frequency Relative frequency Percentage frequency
A 132 132/440 = 0.3 30%
B 88 88/440 = 0.2 20%
C 176 176/440 = 0.4 40%
D 44 44/440 = 0.1 10%
Total 440 1 100%